# Exploring the Performance-Reproducibility Trade-off in Quality-Diversity

### v1

@@ Abstract
Quality-Diversity (QD) algorithms have exhibited promising results across many domains and applications. However, uncertainty in fitness and behaviour estimations of solutions remains a major challenge when QD is used in complex real-world applications. While several approaches have been proposed to improve the performance in uncertain applications, many fail to address a key challenge: determining how to prioritise solutions that perform consistently under uncertainty, in other words, solutions that are reproducible. Most prior methods improve fitness and reproducibility jointly, ignoring the possibility that they could be contradictory objectives. For example, in robotics, solutions may reliably walk at 90% of the maximum velocity in uncertain environments, while solutions that walk faster are also more prone to falling over. As this is a trade-off, neither one of these two solutions is “better” than the other. Thus, algorithms cannot intrinsically select one solution over the other, but can only enforce given preferences over these two contradictory objectives. In this paper, we formalise this problem as the performance-reproducibility trade-off for uncertain QD. We propose four new a-priori QD algorithms that find optimal solutions for given preferences over the trade-offs. We also propose an a-posteriori QD algorithm for when these preferences cannot be defined in advance. Our results show that our approaches successfully find solutions that satisfy given preferences. Importantly, by simply accounting for this trade-off, our approaches perform better than existing uncertain QD methods. This suggests that considering the performance-reproducibility trade-off unlocks important stepping stones that are usually missed when only performance is optimised.
@@ I Introduction
Many problems, such as decision-making and content generation, benefit greatly from a variety of solutions and options. This provides a range of choices of high-performing solutions to the final user [1] or constitutes reservoirs of alternative solutions in unexpected situations [2]. Quality-Diversity (QD) algorithms [3, 4, 5] have attracted attention in recent years as a method to find such collections of diverse and high-performing solutions. QD has been applied to a range of problems; finding diverse gaits in robotics [2], creating innovative designs [6], or generating content for video games [7]. While uncovering diversity remains the primary usage of these algorithms, they have also proven promising in helping to solve deceptive tasks [8, 9], as they discover stepping stones toward more promising solutions that can hardly be found with directed search [10].
@@
  • We develop $4$ new UQD approaches when preferences are available a-priori. Interestingly, one of them constitutes the general case of the previously proposed ME-LS [15].
@@ II-B Solution Reproducibility
Feature reproducibility is important for QD because of two main reasons. First, downstream solutions are usually sampled from final QD archives based on their features [2, 22]. Feature reproducibility ensures downstream users reliably obtaining solutions with the desired features which is critical to the usefulness of the final archive. Second, solutions with poor feature reproducibility might be added to multiple cells during optimisation, causing it to collapse into just a few final solutions with similar features and resulting in significantly reduced diversity, defeating the purpose of QD. For these reasons, while some work study both feature and fitness reproducibility [11], most UQD work consider feature reproducibility only [17, 15]. This work also focuses on feature reproducibility. From here, we refer to it simply as “reproducibility”, as done in previous work.
@@ II-B1 Emergence of reproducibility
II-B2 Distinction from robustness
Reproducibility shares similarities with robustness, a common concept in Evolutionary Algorithms [24]. Reproducibility refers to the ability of a solution to replicate consistently its fitness or features when evaluated multiple times under the same conditions. Robustness refers to the same ability but when evaluated multiple times under different conditions. So reproducibility is concerned with enforcing repeatability under the training distributions, while robustness is concerned with enforcing repeatability under multiple testing distributions. Some approaches to robustness, such as domain randomisation [25], train on large sets of distributions, hoping these sets would cover and encompass the possible testing distributions. In other words, these approaches aim to convert the robustness problem into a reproducibility problem.
@@ II-B3 Quantifying reproducibility and performance
We purposefully define reproducibility as a property of solutions, in an estimator-agnostic fashion. Several estimators have been proposed to quantify the expected performance, such as the mean or median, and to quantify the reproducibility, such as the standard deviation (std) or spread. While a wider overview is given in Appendix, the algorithms proposed in this work can be used with any estimator. In the following, we use mean for the performance and std for the reproducibility.
@@ II-C Uncertain Quality-Diversity
QD algorithms assume the evaluation of solutions return a reliable estimate of their fitness $f$ and feature $d$, as this is required to compare solutions when adding to the archive $\mathcal{A}$. Thus, they do not provide any mechanism to handle cases where these quantities cannot be reliably estimated from a single evaluation. In robotic tasks, for example, stochastic dynamics might lead solutions to perform slightly differently from one evaluation to another. In such environments, QD algorithms tend to keep solutions that obtain “lucky” evaluations, i.e. outliers whose performance or diversity is overestimated. Consider a robotic controller that only walks the robot $1\%$ of the time when conditions are favourable, if it gets lucky during its single evaluation and manages to walk, it would be kept in the QD archive. These “lucky” solutions are then returned instead of truly diverse or good-performing ones, thus limiting the effectiveness of QD. Therefore, Uncertain QD (UQD) [11] design algorithms for uncertain tasks, where the fitness and feature are no longer fixed values but distributions over possible values: $f\sim\mathcal{D}_{f}$ and $d\sim\mathcal{D}_{d}$.
@@ II-C3 Other UQD approaches
Deep-Grid remove the need for sampling by using neighbouring solutions as proxy samples [14]. Alternatively, multiple works have highlighted the benefits of gradient-based mutations for UQD applications [30, 31]. Both these approaches do not easily allow reproducibility to be accounted for so we do not consider them in this work. ARIA [17] is an optimisation module that improves the performance and reproducibility of the solutions contained in the final collection of any QD algorithm. While this approach is close to our work, it can be run on top of another UQD algorithm, making it hard to compare. Additionally, ARIA requires an order of magnitude more evaluations than any UQD algorithm. Thus, we do not include it in our results. We expand on these choices in Appendix.
@@ III Problem definition
$\begin{split}\max_{\mathcal{A}}\left[\sum_{i\in\mathcal{A}}{\mathbb{P}_{f_{i}\sim\mathcal{D}_{f}}\left[f_{i}\right]}\right]\text{ s.t }\forall i\in\mathcal{A},\left[\mathbb{P}_{d_{i}\sim\mathcal{D}_{d}}\left[d_{i}\right]\right]\in\textrm{cell}_{i}\end{split}$ (1)
Where $\mathbb{P}$ denotes any estimator of performance (e.g. mean or median; see Section II-B3) and $\textrm{cell}_{i}$ refers to the feature niche (or cell) of solution $i$ in the QD archive $\mathcal{A}$ (see Section II-A). In other words, the aim of UQD algorithms is to fill each cell of the archive $\mathcal{A}$ with the solution that has highest chance of belonging to a particular cell and maximum estimated performance, according to the estimator $\mathbb{P}$.
@@
$\begin{split}\max_{\mathcal{A}}\left[\sum_{i\in\mathcal{A}}{\mathbb{P}_{f_{i}\sim\mathcal{D}_{f}}\left[f_{i}\right]}\right]&\text{ s.t }\forall i\in\mathcal{A},\left[\mathbb{P}_{d_{i}\sim\mathcal{D}_{d}}\left[d_{i}\right]\right]\in\textrm{cell}_{i}\\ &\text{ s.t }\forall i\in\mathcal{A},\max\left[\mathbb{R}_{d_{i}\sim\mathcal{D}_{d}}\left[d_{i}\right]\right]\end{split}$ (2)
Where $\mathbb{R}$ denotes any estimator of the reproducibility (e.g.std). Even if it is not formulated as such, this new objective (Equation 2) corresponds to the one optimised by algorithms proposed in prior work [17, 15]. However, this paper aim to stress that this objective is, in most cases, unattainable. In many UQD applications, ideal solutions which have both high fitness and high reproducibility do not exist. The fitness and reproducibility of solutions might be orthogonal objectives. For example, in complex robotic control environments, fast-running solutions are intrinsically more brittle than more conservative slower solutions. Thus, optimisation in the UQD setting implies making a trade-off between performance and reproducibility, an issue that has been largely underestimated in previous works.
@@
$\begin{split}\max_{\mathcal{A}}\left[\sum_{i\in\mathcal{A}}{\mathbb{P}_{f_{i}\sim\mathcal{D}_{f}}\left[f_{i}\right]}\right]\text{ s.t }\forall i\in\mathcal{A},\left[\mathbb{P}_{d_{i}\sim\mathcal{D}_{d}}\left[d_{i}\right]\right]\in\textrm{cell}_{i}\\ \text{ s.t }\forall i\in\mathcal{A},\textrm{preference}\left[\mathbb{P}_{f_{i}\sim\mathcal{D}_{f}}\left[f_{i}\right],\mathbb{R}_{d_{i}\sim\mathcal{D}_{d}}\left[d_{i}\right]\right]\\ \end{split}$ (3)
Section II-C defines the two problems commonly identified in UQD. In this work, we propose a third one: the performance-reproducibility problem, that corresponds to enforcing preferences over the trade-off between performance and reproducibility. This last objective (Equation 3) encompass these three problems.
@@ IV-B2 Adaptive-sampling weighted-sum
IV-C A-priori Approaches: Dominance-based
The $\delta$-parametrisation also allows the implementation of a dominance-based archive addition. When choosing if a new solution $i$ should replace an elite $e$, instead of relying on fitness-only, we use the following criteria (also in Figure 3):
@@ IV-C A-priori Approaches: Dominance-based
IV-C1 Fixed-sampling dominance-addition
fixed-sampling approaches (Section II-C1) can integrate dominance-addition by sampling solutions $N$ times to estimate their fitness and reproducibility, and add to the archive based on Equation 5.
Interestingly, the previously introduced ME-LS approach [15] (see Section II-C1) is a particular case of fixed-sampling dominance-addition with $\delta_{f}{}=0$ and $\delta_{r}{}=0$.
IV-C2 Adaptive-sampling dominance-addition
The integration in adaptive-sampling approaches (Section II-C2) raises the question of non-stationarity already mentioned in the weighted-sum case, and hence requires the use of a depth. However, a depth of $d$ would usually keep the $d$ best-performing solutions encountered so far; while such ranking is easy to obtain for single-score comparison, it becomes more complex when using dominance-addition. To overcome this, we propose the following: when a new solution is considered for addition within a cell, it is compared to the elites one by one in descending order until it dominates one of them and replaces it. The replaced elite is then similarly compared to the less-performing elites of the cell.
@@ V-A1 Algorithms
  • ME-Dominance: a-priori dominance fixed-sampling.
@@
  • AS-Dominance: a-priori dominance adaptive-sampling.
@@
Uncertainty Noise on initial joint positions and velocities.
@@ V-A2 Tasks
We consider the robotic control tasks in Table I, used in previous work [11, 30, 15, 17, 37] (further information provided in Appendix). For $\delta_{f}{}$ and $\delta_{r}{}$, we ran some initial experiments to determine the range of fitness and reproducibility values for each task, then, we picked $10\%$ of these ranges for unidirectional tasks and $20\%$ for omnidirectional tasks as values for $\delta_{f}{}$ and $\delta_{r}{}$. We emphasise that we did not optimise these values but picked them arbitrarily to simulate realistic user’s preferences.
@@ V-A3 Metrics
We compare performance using the Corrected QD-Score and the Reproducibility-Score. To calculate the Corrected QD-Score, we approximate the “real” performance of each solution as the median of $512$ reevaluations (shown to be a reasonable number in Flageat et al. [11]) and add all of them back into an empty archive known as the Corrected Archive. The Corrected QD-Score is the QD-Score [3] of this Corrected archive. We also approximate the reproducibility of solutions using these $512$ reevaluations. The Reproducibility-Score is the sum of all the reproducibilities.
@@ V-B Results
Figure 4: Robotic tasks results: (top) Corrected QD-Score, displaying the quality and diversity of the final archive, and (bottom) Reproducibility-Score, quantifying the reproducibility of the final solutions. We report the median and quartiles across $10$ seeds. Each plot is split (by horizontal lines) into two parts: fixed-sampling approaches (both baselines and proposed approaches) and adaptive-sampling (both baselines and proposed approaches).
@@ V-B2 Detailed comparison with UQD baselines
As mentioned earlier, our $3$ fixed-sampling approaches outperform previous fixed-sampling approaches in Corrected QD-Score on Walker and Hexapod. The only exception is the Ant task, known to be difficult to solve with fixed-sampling [11, 30]. We hypothesise that this is because the task is deceptive: policies that walk in a reliable manner are rare and are surrounded in the search space by more brittle solutions that tend to fall easily. Thus, reaching areas of the search space where solutions walk consistently requires using brittle solutions as stepping stones. Due to their fixed number of reevaluations, fixed-sampling approaches systematically reject these intermediate stepping stones. They get quickly stuck with small archives of a few solutions and therefore there is almost no renewal of the pool of parents. On the contrary, adaptive-sampling approaches keep these stepping stone solutions and are therefore able to overcome this deceptive area. Our two new adaptive-sampling approaches outperform previous adaptive-sampling approaches on Walker ($p<0.001$) and perform similarly on Hexapod. However, while our weighted adaptive-sampling approach also outperform baselines on Ant ($p<0.001$), the dominance variant appears to obtain bimodal results and not systematically solve the task. We hypothesis this might be due to the deceptive structure of the task, that only some replications of this approach manage to overcome.
In term of Reproducibility-Score, our approaches outperform Vanilla-ME, ME-Random and ME-Sampling ($p<0.01$). As could be expected, they perform similarly to ME-LS, which is a particular case of the proposed ME-Dominance approach (see Section IV-C). ME-Sampling-Reproducibility, which optimises only for reproducibility, proves better on Walker ($p<0.001$). This is likely due to the value chosen for $\delta_{f}{}$ and $\delta_{r}{}$ that do not allow our approaches to keep the most reproducible solutions of all. Our $2$ AS approaches also outperform Vanilla-AS on Walker ($p<0.001$) and on Ant for AS-Weighted only ($p<0.001$), but perform similarly on Hexapod. As the Hexapod is controlled using an open-loop controller, the reproducibility of solutions is probably capped and it seems likely that all our approaches have reach a plateau that Vanilla-AS was already reaching.

### v2

@@ Abstract
Quality-Diversity (QD) algorithms have exhibited promising results across many domains and applications. However, uncertainty in fitness and behaviour estimations of solutions remains a major challenge when QD is used in complex real-world applications. While several approaches have been proposed to improve the performance in uncertain applications, many fail to address a key challenge: determining how to prioritise solutions that perform consistently under uncertainty, in other words, solutions that are reproducible. Most prior methods improve fitness and reproducibility jointly, ignoring the possibility that they could be contradictory objectives. For example, in robotics, solutions may reliably walk at 90% of the maximum velocity in uncertain environments, while solutions that walk faster are also more prone to falling over. As this is a trade-off, neither one of these two solutions is “better” than the other. Thus, algorithms cannot intrinsically select one solution over the other, but can only enforce given preferences over these two contradictory objectives. In this paper, we formalise this problem as the performance-reproducibility trade-off for uncertain QD. We propose two new a-priori QD algorithms that find efficient solutions for given preferences over the trade-offs. We also propose an a-posteriori QD algorithm for when these preferences cannot be defined in advance. Our results show that our approaches successfully find solutions that satisfy given preferences. Importantly, by simply accounting for this trade-off, our approaches perform better than existing uncertain QD methods. This suggests that considering the performance-reproducibility trade-off unlocks important stepping stones that are usually missed when only performance is optimised.
@@ I Introduction
Many problems, such as decision-making and content generation, benefit greatly from a variety of solutions and options. This provides a range of choices of high-performing solutions to the final user [1] or constitutes reservoirs of alternative solutions in unexpected situations [2]. Quality-Diversity (QD) algorithms [3, 4, 5] have attracted attention in recent years as a method to find such collections of diverse and high-performing solutions. QD has been applied to a range of problems: finding diverse gaits [2] or grasps [6] in robotics, creating innovative designs [7, 8], or generating content for video games [9, 10]. While uncovering diversity remains the primary usage of these algorithms, they have also proven promising in helping to solve deceptive tasks [11, 12], as they discover stepping stones toward more promising solutions that can hardly be found with directed search [13].
@@
  • We develop $2$ new UQD approaches when preferences are available a-priori. Interestingly, one of them constitutes the general case of the previously proposed ME-LS [18].
@@ II-B Solution Reproducibility
Feature reproducibility is important for QD because of two main reasons. First, when considering the final archive produces by a QD algorithm, downstream solutions are usually sampled from this archive based on their features [2, 25]. Feature reproducibility ensures downstream users reliably obtain solutions with the desired features which is critical to the usefulness of the final archive. Second, when considering the QD optimisation process, solutions with poor feature reproducibility are likely to be added to multiple cells during optimisation, causing the archive to collapse into just a few final solutions with similar features and resulting in significantly reduced diversity in the archive, defeating the purpose of QD. For these reasons, while some work study both feature and fitness reproducibility [14], most UQD work consider feature reproducibility only [20, 18]. This work also focuses on feature reproducibility. From here, we refer to it simply as “reproducibility”, as done in previous work.
@@ II-B1 Emergence of reproducibility
II-B2 Link to robustness
Reproducibility shares similarities with robustness, a common concept in Robotics and Evolutionary Algorithms [27]. Reproducibility refers to the ability of a solution to replicate consistently its fitness or features when evaluated multiple times within the same uncertain environment, encountered during training. Robustness, on the other hand, extends this concept to new variations of the uncertain environment that are only encountered during testing. In other words, reproducibility encapsulates whether a solution performs the same for multiple runs of in-distribution environments whereas robustness captures whether a solution can still perform for out-of-distribution ones. For example, in the robotic example of Figure 1, a reproducible solution would be expected to perform consistently during training and testing within the same environment. In contrast, a robust solution would be expected to perform reliably even in a slightly altered environment, such as one with different friction levels, even if these variations were not encountered during training. Approaches such as domain randomisation [28] aim to convert the robustness problem into a reproducibility problem by training on large sets of environment variations, hoping these sets would encompass the possible testing environments.
@@ II-B3 Quantifying reproducibility and performance
We purposefully define reproducibility as a property of solutions, in an estimator-agnostic fashion. Several estimators have been proposed to quantify the expected performance, such as the mean or median, and to quantify the reproducibility, such as the standard deviation (std) or spread. While a wider overview is given in Appendix B, the algorithms proposed in this work can be used with any estimator. In the following, we use mean for the performance and std for the reproducibility.
@@ II-C Uncertain Quality-Diversity
QD algorithms assume the evaluation of solutions returns a reliable estimate of their fitness $f$ and feature $d$, as this is required to compare solutions when adding to the archive $\mathcal{A}$. Thus, they do not provide any mechanism to handle cases where these quantities cannot be reliably estimated from a single evaluation. In robotic tasks, for example, stochastic dynamics might lead solutions to perform slightly differently from one evaluation to another. In such environments, QD algorithms tend to keep solutions that obtain “lucky” evaluations, i.e. outliers whose performance or diversity is overestimated. Consider a robotic controller that only walks the robot $1\%$ of the time when conditions are favourable, if it gets lucky during its single evaluation and manages to walk, it would be kept in the QD archive. These “lucky” solutions are then returned instead of truly diverse or good-performing ones, thus limiting the effectiveness of QD. Therefore, Uncertain QD (UQD) [14] designs algorithms for uncertain tasks, where the fitness and feature are no longer fixed values but distributions over possible values: $f\sim\mathcal{D}_{f}$ and $d\sim\mathcal{D}_{d}$.
@@ II-C3 Other UQD approaches
Deep-Grid remove the need for sampling by using neighbouring solutions as proxy samples [17]. Alternatively, multiple works have highlighted the benefits of gradient-based mutations for UQD applications [33, 34]. Both these approaches do not easily allow reproducibility to be accounted for so we do not consider them in this work. ARIA [20] is an optimisation module that improves the performance and reproducibility of the solutions contained in the final collection of any QD algorithm. While this approach is close to our work, it can be run on top of another UQD algorithm, making it hard to compare. Additionally, ARIA requires an order of magnitude more evaluations than any UQD algorithm. Thus, we do not include it in our results. We expand on these choices in Appendix A.
@@ III Problem definition
$\begin{split}\max_{\mathcal{A}}\left[\sum_{e\in\mathcal{A}}{\mathbb{P}_{f_{e}\sim\mathcal{D}_{f}}\left[f_{e}\right]}\right]\text{ s.t }\forall e\in\mathcal{A},\left[\mathbb{P}_{d_{e}\sim\mathcal{D}_{d}}\left[d_{e}\right]\right]\in\textrm{cell}_{e}\end{split}$ (1)
Where $\mathbb{P}$ denotes any estimator of performance (e.g. mean or median; see Section II-B3, we note that it does not have to be the same for the fitness and features) and $\textrm{cell}_{e}$ refers to the feature niche (or cell) of solution $e$ in the QD archive $\mathcal{A}$ (see Section II-A). In other words, the aim of UQD algorithms is to fill each cell of the archive $\mathcal{A}$ with the solution that has highest chance of belonging to a particular cell and maximum estimated performance, according to the estimator $\mathbb{P}$.
@@
$\begin{split}\max_{\mathcal{A}}\left[\sum_{e\in\mathcal{A}}{\mathbb{P}_{f_{e}\sim\mathcal{D}_{f}}\left[f_{e}\right]}\right]&\text{ s.t }\forall e\in\mathcal{A},\left[\mathbb{P}_{d_{e}\sim\mathcal{D}_{d}}\left[d_{e}\right]\right]\in\textrm{cell}_{e}\\ &\text{ s.t }\forall e\in\mathcal{A},\max\left[\mathbb{R}_{d_{e}\sim\mathcal{D}_{d}}\left[d_{e}\right]\right]\end{split}$ (2)
Where $\mathbb{R}$ denotes any estimator of the reproducibility (e.g. negative std). Even if it is not formulated as such, this new objective (Equation 2) corresponds to the one optimised by algorithms proposed in prior work [20, 18]. However, this paper aim to stress that this objective is, in most cases, unattainable. In many UQD applications, ideal solutions which have both high fitness and high reproducibility do not exist. The fitness and reproducibility of solutions might be orthogonal objectives. For example, in complex robotic control environments, fast-running solutions are intrinsically more brittle than more conservative slower solutions. Thus, optimisation in the UQD setting implies making a trade-off between performance and reproducibility, an issue that has been largely underestimated in previous works.
@@
$\begin{split}\max_{\mathcal{A}}\left[\sum_{e\in\mathcal{A}}{\mathbb{P}_{f_{e}\sim\mathcal{D}_{f}}\left[f_{e}\right]}\right]\text{ s.t }\forall e\in\mathcal{A},\left[\mathbb{P}_{d_{e}\sim\mathcal{D}_{d}}\left[d_{e}\right]\right]\in\textrm{cell}_{e}\\ \text{ s.t }\forall e\in\mathcal{A},\textrm{preference}\left[\mathbb{P}_{f_{e}\sim\mathcal{D}_{f}}\left[f_{e}\right],\mathbb{R}_{d_{e}\sim\mathcal{D}_{d}}\left[d_{e}\right]\right]\\ \end{split}$ (3)
Where preference refers to any mapping given by a final user that specifies which solution should be preferred over another given their respective fitness and reproducibility. Section II-C defines the two problems commonly identified in UQD. In this work, we propose a third one: the performance-reproducibility problem, that corresponds to enforcing preferences over the trade-off between performance and reproducibility. Equation 3 encompasses these three problems.
@@ IV-B2 Adaptive-sampling weighted-sum
IV-C A-priori Approaches: Delta Comparison
The $\delta$-parametrisation also allows the implementation of a delta-comparison-based archive addition. When choosing if a new solution $i$ should replace an elite $e$, instead of relying on fitness-only, we use the following criteria (also in Figure 3):
@@ IV-C A-priori Approaches: Delta Comparison
IV-C1 Fixed-sampling delta-comparison
fixed-sampling approaches (Section II-C1) can integrate delta-comparison by sampling solutions $N$ times to estimate their fitness and reproducibility, and add to the archive based on Equation 5.
Interestingly, the previously introduced ME-LS approach [18] (see Section II-C1) is a particular case of fixed-sampling delta-comparison with $\delta_{f}{}=0$ and $\delta_{r}{}=0$.
IV-C2 Adaptive-sampling delta-comparison
The integration in adaptive-sampling approaches (Section II-C2) raises the question of non-stationarity already mentioned in the weighted-sum case, and hence requires the use of a depth. However, a depth of $d$ would usually keep the $d$ best-performing solutions encountered so far; while such ranking is easy to obtain for single-score comparison, it becomes more complex when using delta-comparison. To overcome this, we propose the following: when a new solution is considered for addition within a cell, it is compared to the elites one by one in descending order until it dominates one of them and replaces it. The replaced elite is then similarly compared to the less-performing elites of the cell.
@@ V-A1 Algorithms
  • ME-Delta: a-priori delta-comparison fixed-sampling.
@@
  • AS-Delta: a-priori delta-comparison adaptive-sampling.
@@
Uncertainty Noise on fitness and feature. Noise on initial joint positions and velocities.
@@ V-A2 Tasks
We consider the robotic control tasks in Table I, used in previous work [14, 33, 18, 20, 39] (further information provided in Appendix C). For $\delta_{f}{}$ and $\delta_{r}{}$, we ran some initial experiments to determine the range of fitness and reproducibility values for each task, then, we picked $10\%$ of these ranges for unidirectional tasks and $20\%$ for omnidirectional tasks as values for $\delta_{f}{}$ and $\delta_{r}{}$. We emphasise that we did not optimise these values but picked them arbitrarily to simulate realistic user’s preferences.
@@ V-A3 Metrics
We compare performance using the Corrected QD-Score and the Reproducibility-Score. To calculate the Corrected QD-Score, we approximate the “real” performance of each solution as the median of $512$ reevaluations (shown to be a reasonable number in Flageat et al. [14]) and add all of them back into an empty archive known as the Corrected Archive. The Corrected QD-Score is the QD-Score [3] of this Corrected archive. We also approximate the reproducibility of solutions using these $512$ reevaluations. To do so, we first get the descriptor variance of the $512$ reevaluations of each solution, normalised within each cell using the maximum observed-variance. This normalisation accounts for the difference in descriptor variance across the descriptor space. The Reproducibility-Score is then computed as the sum over the archive of 1 - normalised variance, to avoid penalising approaches that find more solutions.
@@ V-B Results
Figure 4: Robotic tasks results: (top) Corrected QD-Score, displaying the quality and diversity of the final archive, and (bottom) Reproducibility-Score, quantifying the reproducibility of the solutions in the final archive. For both metrics, higher score is better. The vertical lines show the median across $10$ replications, the boxes the quartiles, the whiskers $1.5$ times the interquartile range, and the dots represent outliers. Each plot is split (by horizontal lines) into two parts: fixed-sampling approaches (both baselines and proposed approaches) and adaptive-sampling (both baselines and proposed approaches).
@@ V-B2 Detailed comparison with UQD baselines
As mentioned earlier, our $3$ fixed-sampling approaches outperform previous fixed-sampling approaches in Corrected QD-Score on Walker and Hexapod. The only exception is the Ant task, known to be difficult to solve with fixed-sampling [14, 33]. We hypothesise that this is because the task is deceptive: policies that walk in a reliable manner are rare and are surrounded in the search space by more brittle solutions that tend to fall easily. Thus, reaching areas of the search space where solutions walk consistently requires using brittle solutions as stepping stones. Due to their fixed number of reevaluations, fixed-sampling approaches systematically reject these intermediate stepping stones. They get quickly stuck with small archives of a few solutions and therefore there is almost no renewal of the pool of parents. On the contrary, adaptive-sampling approaches keep these stepping stone solutions and are therefore able to overcome this deceptive area. Our two new adaptive-sampling approaches outperform previous adaptive-sampling approaches on Walker ($p<0.001$) and perform similarly on Hexapod. However, while our weighted adaptive-sampling approach also outperform baselines on Ant ($p<0.001$), the delta-comparison variant appears to obtain bimodal results and not systematically solve the task. We hypothesise this might be due to the deceptive structure of the task, which only some replications manage to overcome.
In term of Reproducibility-Score, our approaches outperform Vanilla-ME, ME-Random and ME-Sampling ($p<0.01$). As could be expected, they perform similarly to ME-LS, which is a particular case of the proposed ME-Delta approach (see Section IV-C). ME-Sampling-Reproducibility, which optimises only for reproducibility, proves better on Walker ($p<0.001$). This is likely due to the value chosen for $\delta_{f}{}$ and $\delta_{r}{}$ that do not allow our approaches to keep the most reproducible solutions of all. Our $2$ AS approaches also outperform Vanilla-AS on Walker ($p<0.001$) and on Ant for AS-Weighted only ($p<0.001$), but perform similarly on Hexapod. As the Hexapod is controlled using an open-loop controller, the reproducibility of solutions is probably capped and it seems likely that all our approaches have reach a plateau that Vanilla-AS was already reaching.